1) A cross was set up between alpacas of genotypes EeFFggHh and EeFfGgHh

A) What proportion of the offspring are expected to be E- F- G- H- ? (E- means EE or Ee.)

B) What proportion of the offspring are expected to be heterozygous for all four genes?

C) What proportion of the offspring are expected to be eeFfggHH?

D) What proportion of the offspring are expected to be homozygous at all four loci.
Answer #1

Answer:

EeFFggHh X EeFfGgHh ---Parents

Ee x Ee = EE (1/4), Ee (1/2) & ee (1/4)

EE + Ee = E_ (1/4+1/2=3/4)

FFx Ff = FF (1/2) & Ff (1/2) Or F_(1)

gg x Gg = Gg (1/2) & gg (1/2)

Hh x Hh = HH (1/4), Hh (1/2) & hh (1/4)

A) The proportion of the offspring are expected to be E- F- G- H- = ¾ * 1 * ½ * ¾ = 9/32

B) The proportion of the offspring are expected to be heterozygous for all four genes = EeFfGgHh = ½ * ½ * ½ * ½ = 1/16

C) The proportion of the offspring are expected to be eeFfggHH = ¼ * ½ * ½ * ¼ = 1/64

D) The proportion of the offspring are expected to be homozygous at all four loci = 1/64

EE FF gg HH = ¼ * ½ * ½ * ¼ = 1/64

EE FF gg hh = ¼ * ½ * ½ * ¼ = 1/64

ee FF gg HH = ¼ * ½ * ½ * ¼ = 1/64

ee FF gg hh = ¼ * ½ * ½ * ¼ = 1/64